# Unitary Designs of Symmetric Local Random Circuits

### v1

@@ Abstract
This theorem applies to arbitrary symmetries including noncommutative ones as long as the semi-universality condition holds. We note that Eq. (12) means $t<\infty$ when $(\mathcal{V}^{\perp}\cap\mathbb{Z})\backslash\{\bm{0}\}=\varnothing$, which consists with the fact that $\zeta$ is an approximate $(G,R)$-symmetric unitary $t$-design for all $t\in\mathbb{N}$ when $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$ is universal for $\mathcal{U}_{n,G,R}$. We show that this theorem also implies the converse below. We also note that when the representation $R$ is given by $T^{\otimes n}$ with a single qudit representation $T$ of $G$, $\mathcal{V}$ is given by $\{\bm{f}(A\otimes\mathrm{I}^{\otimes n-k})|A\in\mathcal{L}_{k,G,T^{\otimes k}}\}$, where $\mathcal{L}_{k,G,T^{\otimes k}}$ is the set of all linear operators on $k$ qubits commuting with $T^{\otimes k}(g)$ for all $g\in G$. In this case, the dimension of $\mathcal{V}$ is constant with respect to the qudit count $n$, because it is no greater than the dimension of $\mathcal{L}_{k,G,T^{\otimes k}}$.
We explain how we can confirm that $\zeta$ is an approximate unitary $t$-design for all $t\in\mathbb{N}$ if and only if $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ is universal for $\mathcal{U}_{n,G,R}$. The “if” part is trivial because we can apply the same argument in the absence of symmetry. For the converse part, whenever $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ is not universal, we can give a not necessarily tight upper bound $u$ on $t$. We can get such $u$ by $u:=(\bm{m},\bm{c}^{+})$ with some $\bm{c}\in(\mathcal{V}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}$ (see below Theorem 1 of Ref. [42] for the proof of the existence of $\bm{c}$).
We also remark that the (not necessarily tight) upper bound $u$ is useful for the numerical computation of the tight bound, because Eq. (12) is equivalent to
$t<\min\left\{u,\min_{\bm{x}\in(\mathcal{V}^{\perp}\cap\mathcal{F}_{u}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}(\bm{m},\bm{x}^{+})\right\}$ (13)
with $\mathcal{F}_{u}:=\{\bm{x}\in\mathbb{R}^{\Lambda}||x_{\lambda}|<u/m_{\lambda}\ \forall\lambda\in\Lambda\}$, and the r.h.s. of Eq. (13) can be computed by enumeration. For the proof of Eq. (13), it is sufficient to show that $\min_{\bm{x}\in\overline{\mathcal{F}_{u}}\cap\mathbb{Z}^{\Lambda}}(\bm{m},\bm{x}^{+})\geq u$ by noting that $[(\mathcal{V}^{\perp}\cap\mathcal{F}_{u}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}]\cup(\overline{\mathcal{F}_{u}}\cap\mathbb{Z}^{\Lambda})\supset(\mathcal{V}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}$. For arbitrary $\bm{x}\in\overline{\mathcal{F}_{u}}\cap\mathbb{Z}^{\Lambda}$, we can take some $\lambda\in\Lambda$ such that $|x_{\lambda}|\geq u/m_{\lambda}$. We can suppose that $x_{\lambda}>0$ without loss of generality, since $\bm{x}\in\overline{\mathcal{F}_{u}}\cap\mathbb{Z}^{\Lambda}$ implies $-\bm{x}\in\overline{\mathcal{F}_{u}}\cap\mathbb{Z}^{\Lambda}$. Then, we get $(\bm{m},\bm{x}^{+})\geq m_{\lambda}x_{\lambda}\geq u$.
@@
Although we present the complete proof of Theorem 2 in Ref. [42], we briefly explain the proof idea. First, we give an intuitive explanation. Since we consider the case when the set of symmetric local gates is semi-universal, the expressibility of symmetric local gate set and that of symmetric global gate set are different only in terms of the relative phases. Therefore, the notion of approximate unitary $t$-design is related to the controllability of the $t$-fold relative phase Such condition is equivalently rephrased as the condition for a certain set of equations having no nontrivial integer solution, which can also be expressed as integer optimization.
Next, we comment on the technical perspective of the proof. By a standard argument, the distribution $\zeta$ is an approximate $(G,R)$-symmetric $t$-design if and only if the commutants of $\Omega_{t}(\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma})$ and $\Omega_{t}(\mathcal{U}_{n,G,R})$ are identical, where $\Omega_{t}$ is the $t$-fold channel defined by $\Omega_{t}(U):=U^{\otimes t}$. Thus for the proof of the “if” part, we prove that every element in $\Omega_{t}(\mathcal{U}_{n,G,R})$ can be generated by $\bigcup_{\gamma\in\Gamma}\Omega_{t}({\mathcal{U}_{n,G,R}^{\gamma}})$ in the sense of algebra. Since the semi-universality is assumed, it is sufficient to show that every element in $\Omega_{t}(Z(\mathcal{U}_{n,G,R}))$ can be generated. For that purpose, we take a certain map from a set of real numbers to $\braket{\Omega_{t}(\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma})}$ and conduct an integration like the Fourier transformation. For the proof of the “only if” part, we prove that we can construct an operator that commutes with $\bigcup_{\gamma\in\Gamma}\Omega_{t}(\mathcal{U}_{n,G,R}^{\gamma})$ but not with $\Omega_{t}(\mathcal{U}_{n,G,R})$. For the construction, we use the totally antisymmetric state defined as |$\ket{\chi(\mathbb{C}^{m})}:=\sum_{\sigma\in\mathfrak{S}_{m}}\mathrm{sgn}(\sigma)\bigotimes_{j=1}^{m}\ket{\sigma(j)}/\sqrt{m!}$ with the symmetric group $\mathfrak{S}_{m}$ of degree $m$. This state is useful because it satisfies $A^{\otimes m}\ket{\chi(\mathbb{C}^{m})}=\mathrm{det}(A)\ket{\chi(\mathbb{C}^{m})}$ for all operators $A$ on $\mathbb{C}^{m}$.
Derivation of Theorem 1 from Theorem 2.— We explain the general procedure to obtain the explicit tight bound on $t$ for a given symmetry such as Theorem 1 by using the general result of Theorem 2. This process consists of three steps. In the first step, we find the concrete expression of $m_{\lambda}$ for the symmetry, and take the basis $\{\bm{v}_{j}\}_{j=1}^{J}$ of $\mathcal{V}$. The multiplicity $m_{\lambda}$ can be obtained by using the representation theory, as we have mentioned just above Theorem 2. As for taking the basis of $\mathcal{V}$, it is important to know the dimension $J$. We can show that $J=1$, $k+1$, and $\lfloor k/2\rfloor+1$ in the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ cases, respectively.
In the second step, we get a not-necessarily tight upper bound $u^{\prime}$ on $t$ by taking a certain $\bm{x}\in(\mathcal{V}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}$. We consider the case when the size of $\Lambda$ is larger than the dimension $J$ of $\mathcal{V}$, because otherwise the gate set $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$ is universal, and thus $\zeta$ is an approximate $(G,R)$-symmetric unitary $t$-design for all $t$. We label the elements of $\Lambda$ as $\lambda_{1}$, $\lambda_{2}$, …, $\lambda_{\#\Lambda}$ such that $m_{\lambda_{1}}\leq m_{\lambda_{2}}\leq\cdots\leq m_{\lambda_{\#\Lambda}}$ and restrict the region of $\bm{x}$ to $\mathcal{D}:=\{\bm{x}\in\mathbb{R}^{\Lambda}|x_{\lambda_{j}}=0\ \forall j\geq J+2\}$. Then, $\mathcal{V}^{\perp}\cap\mathcal{D}$ is one-dimensional, and we can take $\bm{p}\in\mathbb{Z}^{\Lambda}$ such that $\mathcal{V}^{\perp}\cap\mathcal{D}\cap\mathbb{Z}^{\Lambda}=\{a\bm{p}|a\in\mathbb{Z}\}$. We define $u^{\prime}:=(\bm{m},\bm{p}^{+})$, which gives an upper bound on $t$.
In the third step, we show that $u^{\prime}$ is the tight bound under the assumption that $u^{\prime}\leq m_{\lambda_{j}}$ for all $j\geq J+2$, which holds for sufficiently large $n$ in the $\mathrm{U}(1)$ and $\mathrm{SU}(2)$ cases. By applying Eq. (13), it is sufficient to show that $\min_{\bm{x}\in(\mathcal{V}^{\perp}\cap\mathcal{F}_{u^{\prime}}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}(\bm{m},\bm{x}^{+})\geq u^{\prime}$. By using the assumption above, we get the inclusion relation $\mathcal{F}_{u^{\prime}}\cap\mathbb{Z}^{\Lambda}\subset\mathcal{D}\cap\mathbb{Z}^{\Lambda}$. We thus have $\min_{\bm{x}\in(\mathcal{V}^{\perp}\cap\mathcal{F}_{u}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}(\bm{m},\bm{x}^{+})\geq\min_{\bm{x}\in(\mathcal{V}^{\perp}\cap\mathcal{D}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}(\bm{m},\bm{x}^{+})=u^{\prime}$.
@@
We illustrate these steps in the $\mathrm{U}(1)$-symmetric and $2$-local case. First, as we have mentioned just above Theorem 2, we have $\Lambda=\{0,1,...,n\}$, and the multiplicity $m_{\lambda}$ is given by $\binom{n}{\lambda}$. In order to take the basis of $\mathcal{V}$, we note that every U(1)-symmetric 2-local operator $A\in\mathcal{L}_{2,G,T^{\otimes 2}}$ can be written in a block diagonal form $A_{0}\oplus A_{1}\oplus A_{2}$ with $A_{j}$’s acting on the eigenspaces of $\mathrm{Z}\otimes\mathrm{I}+\mathrm{I}\otimes\mathrm{Z}$ with eigenvalues $2-2j$, respectively. Then, $f_{\lambda}(A\otimes\mathrm{I}^{\otimes n-2})$ is given by $\sum_{j=0}^{2}\mathrm{tr}(A_{j})\binom{n-2}{\lambda-j}$, which implies that we can take $\bm{v}_{j}=(\binom{n-2}{\lambda-j})_{\lambda\in\Lambda}$ with $j=0,1,2$ as a basis of $\mathcal{V}$, where we define $\binom{a}{b}:=0$ when $b\not\in\{0,1,...,a\}$. Next, we define $\mathcal{D}:=\{\bm{x}\in\mathbb{R}^{\Lambda}|x_{\lambda}=0\ \forall\lambda\in\{2,3,...,n-2\}\}$. Then, $\mathcal{V}^{\perp}\cap\mathcal{D}\cap\mathbb{Z}^{\Lambda}$ can be expressed as $\{a\bm{p}|a\in\mathbb{Z}\}$ with $\bm{p}\in\mathbb{Z}^{\Lambda}$ defined by $(p_{0},p_{1},p_{n-1},p_{n})=(1,-(n-2),n-2,-1)$ and $p_{\lambda}=0$ for all $\lambda\in\{2,3,...,n-2\}$. We thus define $u^{\prime}:=2(n-1)$, which is an upper bound on $t$. Finally, we note that the assumption holds for all $n\geq 3$, i.e. $u^{\prime}=2(n-1)\leq\binom{n}{2}\leq m_{\lambda_{j}}$ for all $j\geq 5$, where we note that there does not exist an element $\lambda_{j}$ with $j\geq 5$ when $n=3$. We therefore conclude that $u^{\prime}=2(n-1)$ is the tight bound on $t$ for all $n$. We can deal with general $k$-local cases with a similar method.
@@
Discussion.— In this work, we presented a formula for finding the maximal order of unitary design generated with symmetric local random circuits for general symmetries and general localities in Theorem 2. By applying this result to the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries, we have explicitly shown the maximal order of unitary designs for general localities in Theorem 1. Although we have focused on a continuous gate set $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$, our results are also applicable to random circuits with an arbitrary gate set as long as the gate set can generate the continuous unitary ensemble $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$. Since Theorem 2 is given in a general form, we can derive the results for other symmetries and localities.
As a future research direction, it is important to investigate the convergence speed of the approximate unitary designs on symmetric and local circuits. Moreover, the conserved quantity in $t$ copies of a system under symmetric local unitaries, which we have constructed in the proof of the “only if” part of Theorem 2, may lead to the discovery of some limitation in the process of thermalization under symmetric local Hamiltonians.
@@
Acknowledgements.— The authors wish to thank Iman Marvian, Hiroyasu Tajima, and Janek Denzler for insightful discussions. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. R.S. is supported by the BMBF (PhoQuan, Grant No. 13N16103). This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S. N.Y. wishes to thank JST PRESTO No. JPMJPR2119, JST ASPIRE Grant Number JPMJAP2316, and the support from IBM Quantum. This work was supported by JST Grant Number JPMJPF2221, JST ERATO Grant Number JPMJER2302, and JST CREST Grant Number JPMJCR23I4, Japan.
@@
Note Added: During the preparation of this article, we became aware of independent work by Austin Hulse, Hanqing Liu, and Iman Marvian, which studies similar questions and was posted on arXiv concurrently with the present paper.

### v3

@@ Abstract
This theorem applies to arbitrary symmetries including noncommutative ones as long as the semi-universality condition holds. We note that when the representation $R$ is given by $T^{\otimes n}$ with a single-qudit representation $T$ of $G$, for example in Eqs. (1), (2), and (3), $\mathcal{C}$ can be simplified to $\{\bm{f}(A\otimes\mathrm{I}^{\otimes n-k})|A\in\mathcal{L}_{k,G,T^{\otimes k}}\}$, where $\mathcal{L}_{k,G,T^{\otimes k}}$ is the set of all linear operators on $k$ qubits commuting with $T^{\otimes k}(g)$ for all $g\in G$. In this case, the dimension of $\mathcal{C}$ is constant with respect to the qudit count $n$, because it is no greater than the dimension of $\mathcal{L}_{k,G,T^{\otimes k}}$.
Two remarks are in order. First, by Theorem 2, we can confirm that $\zeta$ is an asymptotic unitary $t$-design for all $t\in\mathbb{N}$ if and only if $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ is universal for $\mathcal{U}_{n,G,R}$. Concretely, when the gate set is universal, it follows that $\mathcal{C}^{\perp}=\{\bm{0}\}$, and thus Eq. (13) means $t<\infty$, where we define the minimization over the empty set as $\infty$. On the other hand, when the gate set is not universal, we can construct an upper bound $t_{0}$ on the r.h.s. of Eq. (13) as follows: We take some $\bm{d}\in(\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}$, and by using $\bm{d}^{+}:=((|d_{\lambda}|+d_{\lambda})/2)_{\lambda\in\Lambda}$, define $t_{0}:=\braket{\bm{m},\bm{d}^{+}}$, which is an upper bound (see Lemma 14 in Ref. [74] for the proof of the existence of such $\bm{d}$). Then, $\zeta$ is not an asymptotic unitary $t_{0}$-design.
Second, the r.h.s. of Eq. (13) can be calculated by an enumerative method by using the upper bound $t_{0}$ mentioned above. Concretely, we take every $\bm{x}\in(\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}$ satisfying $|x_{\lambda}|<t_{0}/m_{\lambda}$ for all $\lambda\in\Lambda$, and check whether there exists $\bm{x}$ such that $\braket{\bm{m},\bm{x}^{+}}<t_{0}$ by enumeration. If there exists such $\bm{x}$, the r.h.s. of Eq. (13) is given by the minimum of $\braket{\bm{m},\bm{x}^{+}}$, and otherwise, it is given by $t_{0}$. This method works because for any $\bm{x}$ satisfying $|x_{\lambda}|\geq t_{0}/m_{\lambda}$ with some $\lambda\in\Lambda$, we have $\braket{\bm{m},\bm{x}^{+}}=\max\{\braket{\bm{m},(\pm\bm{x})^{+}}\}\geq\max\{m_{\lambda}(\pm x_{\lambda})^{+}\}=m_{\lambda}|x_{\lambda}|\geq t_{0}$. We note that there are some cases when the enumeration is unnecessary, such as $\mathrm{U}(1)$-symmetric case for sufficiently large $n$, which is explained in the proof below.
@@
We find that it is instructive to present the intuitive proof idea of Theorem 2 (see Theorem 1 of Ref. [74] for the complete proof). The notion of asymptotic unitary design is directly connected to the expressibility of the available gate sets. Since we assume semi-universality, the difference in the expressibility of symmetric local/global unitaries appears only in terms of the relative phases. Therefore, the condition of asymptotic $t$-design reduces to the possibility of estimating the component of the relative phases for any sum of $t$ relative phases. This condition is equivalently rephrased as the condition for a certain set of equations having no nontrivial integer solution, which can also be expressed as integer optimization.
Derivation of Theorem 1 from Theorem 2.— Although we guide readers to Theorems 2, 3, and 5 in Ref. [74] for rigorous proofs, we explain the general procedure to get explicit values of the r.h.s. of Eq. (13) under appropriate conditions on $n$ in three steps. Note that the latter two steps are in similar to the numerical method mentioned above.
In the first step, we get the explicit expressions of $\bm{m}$ and the basis of $\mathcal{C}$ by using the representation theory. We focus on the case when the size of $\Lambda$ is larger than the dimension $J$ of $\mathcal{C}$, since Eq. (13) trivially means $t<\infty$ otherwise. In the second step, we take an upper bound $t_{0}$ on r.h.s. of Eq. (13) as follows: we define $\Lambda^{\prime}$ as the set of labels $\lambda$ corresponding to $J+1$ smallest $m_{\lambda}$’s, and define the $(J+1)$-dimensional space $\mathcal{D}:=\{\bm{x}\in\mathbb{R}^{\Lambda}|x_{\lambda}=0\ \forall\lambda\in\Lambda\backslash\Lambda^{\prime}\}$. Then, $\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda}\cap\mathcal{D}$ is a $1$-dimensional lattice. We take a nonzero lattice point $\bm{d}$ closest to $\bm{0}$, and define $t_{0}:=\braket{\bm{m},\bm{d}^{+}}$. In the third step, we show that the r.h.s. is exactly given by $t_{0}$ if $m_{\lambda}\geq t_{0}$ for all $\lambda\in\Lambda\backslash\Lambda^{\prime}$, which holds for sufficiently large $n$ in the $\mathrm{U}(1)$ and $\mathrm{SU}(2)$ cases. This can be confirmed by noting that if we take arbitrary $\bm{x}\in(\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\mathcal{D}$, then we have $\braket{\bm{m},\bm{x}^{+}}\geq m_{\lambda}|x_{\lambda}|\geq t_{0}$.
@@
We illustrate this procedure by taking the $\mathrm{U}(1)$-symmetric and $2$-local case as the simplest nontrivial example. First, as we have mentioned just above Theorem 2, we have $\Lambda=\{0,1,...,n\}$, and the multiplicity $m_{\lambda}$ is given by $\binom{n}{\lambda}$. In order to take the basis of $\mathcal{C}$, we note that every $A\in\mathcal{L}_{2,G,T^{\otimes 2}}$ can be written in a block diagonal form $A_{0}\oplus A_{1}\oplus A_{2}$ with $A_{j}$’s acting on the eigenspaces of $\mathrm{Z}\otimes\mathrm{I}+\mathrm{I}\otimes\mathrm{Z}$ with eigenvalues $2-2j$, respectively. Thus, $f_{\lambda}(A\otimes\mathrm{I}^{\otimes n-2})$ is given by $\sum_{j=0}^{2}\mathrm{tr}(A_{j})\binom{n-2}{\lambda-j}$, which implies that we can take $\bm{c}_{j}=(\binom{n-2}{\lambda-j})_{\lambda\in\Lambda}$ with $j=0,1,2$ as a basis of $\mathcal{C}$, where we define $\binom{a}{b}:=0$ when $b\not\in\{0,1,...,a\}$. Second, we take an upper bound on the r.h.s. of Eq. (13). Since $\mathcal{C}$ is $3$-dimensional, we define $\Lambda^{\prime}:=\{0,1,n-1,n\}$ with its size $4$. Then, the set of $\bm{x}\in\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda}$ satisfying $x_{\lambda}=0$ for all $\lambda\in\Lambda\backslash\Lambda^{\prime}$ forms a $1$-dimensional lattice $\{a\bm{d}|a\in\mathbb{Z}\}$ with $\bm{d}:=(1,-(n-2),0,...,0,n-2,-1)$. We define $t_{0}:=\braket{\bm{m},\bm{d}^{+}}=2(n-1)$. Third, we show that it equals the r.h.s. of Eq. (13) when $n\geq 4$. For the proof of this, it is sufficient to confirm that when $\bm{x}$ is not in the $1$-dimensional lattice mentioned above, we can take some $\lambda\in\Lambda\backslash\Lambda^{\prime}$ satisfying $x_{\lambda}\neq 0$, and we have $\braket{\bm{m},\bm{x}^{+}}\geq m_{\lambda}|x_{\lambda}|\geq m_{\lambda}\geq\binom{n}{2}\geq 2(n-1)=t_{0}$. We can use a similar method for general locality. We present the proof for the $\mathbb{Z}_{2}$ and $\mathrm{SU}(2)$ symmetries in Appendix A.
@@
Discussion.— In this work, we have presented a formula for finding the maximal order of unitary design generated with symmetric local random circuits for general symmetries and general localities in Theorem 2. By applying this result to the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries, we have explicitly shown the maximal order of unitary designs for general localities in Theorem 1. Although we have focused on a continuous gate set $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$, our results are also applicable to random circuits with arbitrary gate sets as long as they generate the continuous unitary ensemble $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$. Since Theorem 2 is given in a general form, it can be used to derive the results for other symmetries.
Our result can serve as a tool for distinguishing whether we can use some gate set for the symmetric versions of quantum information processing tasks, which may, for example, lead to extending the symmetric classical shadow [77, 78] to other symmetries and gate sets. We note that symmetric Clifford gates do not necessarily generate symmetric unitary $3$-designs or even $2$-designs [71], but instead we may be able to use symmetric local random gates to generate symmetric unitary designs. Moreover, since our results hold for any order of unitary design, it would also be useful for investigating the symmetric versions of notions involving $4$th or higher-order unitary designs, such as higher-order randomized benchmarking [79] and out-of-time-order higher point correlators [47].
As future research directions, it is important to investigate the convergence speed of the asymptotic unitary designs on symmetric local random circuits. Additionally, it would be interesting to investigate the effect of our result on entanglement growth [80] and deep thermalization [81] under symmetric local random unitary dynamics. Clarifying the constraints deriving from the existence of the upper bound on the order of unitary designs will enable us to gain a deeper understanding of the many-body dynamics under Hamiltonians composed of symmetric local interactions.
@@
Acknowledgements.— The authors wish to thank Iman Marvian, Hiroyasu Tajima, Janek Denzler, and Zongping Gong for insightful discussions. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. R.S. is supported by the BMBF (PhoQuant, Grant No. 13N16103). This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S. N.Y. wishes to thank JST PRESTO No. JPMJPR2119, JST ASPIRE Grant Number JPMJAP2316, and the support from IBM Quantum. This work was supported by JST Grant Number JPMJPF2221, JST ERATO Grant Number JPMJER2302, and JST CREST Grant Number JPMJCR23I4, Japan.
@@
Note Added: During the preparation of this article, we became aware of independent work by Austin Hulse, Hanqing Liu, and Iman Marvian  [82], which studies similar questions and was posted on arXiv concurrently with the present paper. Both have arrived at the same result on the maximal order of unitary designs under the $\mathrm{U}(1)$ and $\mathrm{SU}(2)$ symmetries. Reference [82] has assumed conjectures about combinatorial identities, which are introduced as Eqs. (86) and (120) of the version 1 of their manuscript for the proof of general $k$-local cases. In our work, we have provided a proof that is independent of any conjectures.